OpenStudy (anonymous):

i hate these and they dont seem to click ! please help. x-3y=5 4x-3y=2

OpenStudy (anonymous):

x - 3y =5 + 3y x = 3y + 5 Put that into the other equation and solve for y: 4(3y + 5) - 3y = 2 Then put the y-value into one of the original equations and solve for x.

OpenStudy (unklerhaukus):

solve the first equation for x, then substitute this value for x into the second equation, and solve for y, use y in either of the original equations to find x